A simple pendulum takes 23.5 seconds to complete 20 cycles. What is the period (T) and frequency of the pendulum?

1
Expert's answer
2021-11-24T12:02:04-0500

"T=\frac{t}{N}=\frac{23.5}{20}=1.175s"

"\nu=\frac{1}{T}=\frac{1}{1.175}=0.851"